in january 2011, the currency was $57 billion and banks held $4 billion of notes and coins, personal chequable deposit were $204, non- personal chequable deposit were $264 , personal non-chequable deposit were $159, non- personal non- chequable deposit were $29 and fixed term deposit were $306.
Q. what was the value of M1 And M2?
Answer #1

M1 = $(57+204+264) billion = $525 billion

M2 = M1 + $(159+29+306) billion = $(525 +159+29+306) billion = $1019 billion.